# On The Hidden Biases of Flow Matching Samplers

### v1

@@ Abstract
We study the implicit bias of flow matching (FM) samplers via the lens of empirical flow matching. Although population FM may produce gradient-field velocities resembling optimal transport (OT), we show that the empirical FM minimizer is almost never a gradient field, even when each conditional flow is. Consequently, empirical FM is intrinsically energetically suboptimal. In view of this, we analyze the kinetic energy of generated samples. With Gaussian sources, both instantaneous and integrated kinetic energies exhibit exponential concentration, while heavy-tailed sources lead to polynomial tails. These behaviors are governed primarily by the choice of source distribution rather than the data. Overall, these notes provide a concise mathematical account of the structural and energetic biases arising in empirical FM.
@@ 1 Introduction
From now on, we assume that all the probability distributions/measures (except the empirical distribution) of the random variables considered are absolutely continuous (i.e., they have densities with respect to the Lebesgue measure), in which case we shall abuse the notation and use the same symbol to denote both the distribution and the density. To maintain the flow of the main text, we defer discussion of related work and all proofs of the theoretical results to the appendix.
@@ 2 Flow Matching (FM) and Conditional Flow Matching (CFM)
Let $p_{0}$ be the source distribution and $p_{1}$ the target distribution (e.g., the data distribution $p^{*}$ or a smoothed version) on $\mathbb{R}^{d}$. We say that $T$ is a transport map if $Z\sim p_{0}$ implies that $T(Z)\sim p_{1}$, in which case we write $T_{\#}p_{0}=p_{1}$, and there exist many such maps. A common generative modeling paradigm aims to learn a transport map $T$ from $p_{0}$ to $p_{1}$ on $\mathbb{R}^{d}$ using $N$ i.i.d. samples $x^{(i)}\sim p_{1}$, where $p_{1}$ is typically unknown. One popular approach under this paradigm is flow matching (FM).
@@
which corresponds to the conditional velocity field $v(t,z|X=x_{1})=\frac{x_{1}-z}{1-t}$. This conditional probability path realizes linear interpolating paths of the form $Z_{t}=(1-t)x_{0}+tx_{1}$ between a (reference) Gaussian sample $x_{0}$ and a data sample $x_{1}$. In practice, regularized versions of rectified flow are preferred for numerical stability (since $v$ blows up as $t\to 1$). A simple version is to modify the conditional probability path to $p_{t}(\cdot|X=x_{1})=\mathcal{N}(tx_{1},(1-(1-\sigma_{min})t)^{2}I_{d})$ for some small $\sigma_{min}>0$, which corresponds to the regularized conditional velocity field $v(t,z|X=x_{1})=\frac{x_{1}-(1-\sigma_{min})z}{1-(1-\sigma_{min})t}$. Another version is to consider a smoothed version of the data distribution $p^{*}$; e.g., $p_{1}=p^{*}\star\mathcal{N}(0,\sigma_{min}^{2}I_{d})$, where $\star$ denotes convolution.
@@ 3 Empirical Flow Matching
where $w_{i}(t,z)=\text{softmax}_{i}\left(\left(-\frac{1}{2(1-t)^{2}}\|z-tx^{(j)}\|^{2}\|\right)_{j\in[N]}\right)$, with $\text{softmax}_{i}$ denoting the $i$th component of the vector obtained after applying the softmax operation. This optimal velocity field is thus a time-dependent weighted average of the $N$ different directions towards the $x^{(i)}$. Similar formula can also be obtained for regularized versions of rectified flow.
@@ 4 FM Through the Lens of Energetics
Hence, the optimization problem (Benamou-Brenier formulation) can be written in its Eulerian form, and minimizes the total kinetic energy:
@@
$\frac{\partial\hat{p}^{*}_{t}}{\partial t}+\nabla\cdot(\hat{v}^{*}\hat{p}^{*}_{t})=0.$ (19)
Remark 1.
By the Helmholtz-Hodge decomposition, any sufficiently smooth vector field in $\mathbb{R}^{d}$ (decaying sufficiently fast at infinity) can be uniquely decomposed into the sum of a gradient field (irrotational) and a divergence-free field (solenoidal) [bhatia2012helmholtz]:
$\hat{v}^{*}(t,z)=\underbrace{-\nabla\Phi(t,z)}_{\text{gradient field}}+\underbrace{u(t,z)}_{\text{divergence-free}},$
for some scalar potential function $\Phi$ and some $u$ such that $\nabla\cdot u=0$. This is analogous to the fact that any matrix can be uniquely decomposed into a symmetric and an antisymmetric part. In the context of OT, the optimality condition requires the velocity to be a gradient field ($u=0$). The presence of a non-zero divergence-free component $u$ implies that the mass transport includes rotational or ”looping” movements that increase the kinetic energy cost without contributing to the net displacement of mass.
@@
Let the empirical target distribution be $\hat{p}_{1}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{(i)}}$. Consider the family of empirical affine flows defined by the conditional probability paths $p_{t}(z|x^{(i)})$ and their corresponding conditional velocity fields $v_{i}(t,z):=v(t,z|x^{(i)})=a_{t}(x^{(i)})z+b_{t}(x^{(i)})$ from Proposition 1. The minimizer $\hat{v}^{*}(t,z)$ of the empirical FM/CFM objective is not a gradient field if and only if the following data-dependent condition holds for all $t\in[0,1]$:
@@
Neural network trained CFM models are implicitly optimizing for a sampler that is not only energetically inefficient (not minimizing the kinetic energy), but also leads to memorization (producing samples that are close enough to those from the training set).
@@
It turns out that we can also obtain a similar bound for the empirical RF model $\hat{v}^{*}(t,z)$ (which is nonlinear in $z$) under the same assumption for $p_{1}$, despite the fact that it does not give rise to an OT map.
@@
Combining these two interpretations, we see that neural network trained CFM models are implicitly optimizing for a sampler that is suboptimal in expected kinetic energy (using more energy on average than the OT solution due to lack of gradient structure), and also leads to memorization. However, regardless of optimality, both OT and CFM models generate samples with exponentially concentrated kinetic energy around typical values, with high-energy samples (which are rare) predominantly found in low-density tail regions.
@@ 5 Conclusion
In these notes, we show that empirical FM optimizes for velocity fields that generally lack a gradient structure and therefore cannot reproduce OT maps or minimize kinetic energy. This explains both the energetic inefficiency and the memorization tendencies observed in practice. Despite this, the generated kinetic energy concentrates sharply: exponentially for Gaussian sources and polynomially for heavy-tailed ones. This shows that the tail behavior of FM samplers is governed mainly by the source distribution, not the underlying data. Understanding these structural biases clarifies the limitations of empirical FM and may guide improved sampler designs, which we leave for future work.
@@ Appendix A Related Work
Flow Matching and Related Models. Flow Matching (FM) and Conditional Flow Matching (CFM) have been developed as scalable alternatives to diffusion-based generative models. Recent work has analyzed their statistical, geometric, and algorithmic foundations, including distributional properties of FM [kunkel2025distribution], particle and bridge-based interpretations [bamberger2025carr], and geometric structure and gauge freedom in learned flows [wanelucidating, horvat2024gauge]. Extensions include guided generation [feng2025guidance], statistical efficiency analyses [mena2025statistical], and rigorous comparisons between FM and optimal transport [hertrich2025relation, wald2025flow]. The kinetic and energetic behavior of flow-based samplers has also been examined in [shaul2023kinetic].
Memorization vs. Generalization in Generative Models. A large body of recent work studies memorization, generalization, and interpolation phenomena in modern generative models. For diffusion models, prior work has analyzed identifiability, overfitting, and deterministic sampling behavior [pidstrigach2022score, yoon2023deterministic]. Further studies provide theoretical and empirical characterizations of interpolation, dataset coverage, and memorization tendencies [lyu2025resolving, scarvelis2023closed, baptista2025memorization, bertrand2025closed, chen2025interpolation]. Broader perspectives on generalization in flow-based and likelihood-based models appear in [zhang2025understanding, gao2024flow, li2024good, ye2025provable, reu2025gradient, li2025generation]. Our work contributes to this line by showing that empirical FM induces a structural bias—non-conservative velocities—that naturally leads to memorization-like behavior even without model approximation error.
@@ B.3 Proof of Proposition 3
Compute the integral explicitly, complete the square, and simplify. ∎
@@
Substituting this into the energy definition:
$E(Y)=\|Y-R^{-1}(Y)\|^{2}$
Since $R^{-1}(R(X))=X$ by definition of the inverse, this simplifies to:
$E=\|R(X)-X\|^{2}.$
Using the definition of $R(X)$:
@@
Let $A=\Sigma_{1}^{1/2}-I_{d}$. Note that $A$ is symmetric and we can consider the eigen-decomposition $A=UDU^{T}$, where $U$ is orthogonal and $D$ is diagonal with elements $d_{i}$. The eigenvalues of $\Sigma_{1}^{1/2}$ are $\sqrt{\lambda_{i}(\Sigma_{1})}$. Thus, the eigenvalues of $A$ are:
$d_{i}=\sqrt{\lambda_{i}(\Sigma_{1})}-1.$
The kinetic energy can then be written as:
@@ B.5 Proof of Theorem 2
For $u\geq U_{t}$ large enough so that $s^{2}=\frac{u}{C_{K}(T)}-1\geq\frac{u}{2C_{K}(T)}$, we have
@@
where $C_{\alpha}^{\prime}:=C_{\alpha}(2C_{K}(T))^{\alpha/2}$. This proves the first inequality in (2) with $\gamma=\alpha/2$.

### v2

@@ Abstract
We study the implicit bias of flow matching (FM) samplers via the lens of empirical flow matching. Although population FM may produce gradient-field velocities resembling optimal transport (OT), we show that the empirical FM minimizer is generally not a gradient field, even when each conditional flow is. Consequently, empirical FM is intrinsically not OT-optimal in the Benamou-Brenier sense. In view of this, we analyze the kinetic energy of generated samples. With Gaussian sources, both instantaneous and integrated kinetic energies exhibit exponential concentration, while heavy-tailed sources lead to polynomial tails. These behaviors are governed primarily by the choice of source distribution rather than the data. Overall, these notes provide a concise mathematical account of the structural and energetic biases arising in empirical FM.
@@ 1 Introduction
From now on, to keep the mathematical machinery light, we assume that all probability distributions/measures of the random variables considered (except the empirical distribution itself) are absolutely continuous, i.e., they admit densities with respect to the Lebesgue measure. We shall abuse notation and use the same symbol to denote both a distribution and its density. To maintain the flow of the main text, we defer the discussion of related work and all proofs of the theoretical results to the appendix.
@@ 2 Flow Matching (FM) and Conditional Flow Matching (CFM)
Let $p_{0}$ be the source distribution and $p_{1}$ the target distribution (e.g., the data distribution $p^{*}$ or a smoothed version) on $\mathbb{R}^{d}$. We say that $T$ is a transport map if $Z\sim p_{0}$ implies that $T(Z)\sim p_{1}$, in which case we write $T_{\#}p_{0}=p_{1}$, and there exist many such maps. A common generative modeling paradigm aims to learn a transport map $T$ from $p_{0}$ to $p_{1}$ on $\mathbb{R}^{d}$ using $N$ i.i.d. samples $x^{(i)}\sim p_{1}$, where $p_{1}$ is typically unknown [peyre2025optimal]. One popular approach under this paradigm is flow matching (FM).
@@
which corresponds to the conditional velocity field $v(t,z|X=x_{1})=\frac{x_{1}-z}{1-t}$. This conditional probability path realizes linear interpolating paths of the form $Z_{t}=(1-t)x_{0}+tx_{1}$ between a (reference) Gaussian sample $x_{0}$ and a data sample $x_{1}$. In practice, regularized versions of rectified flow are preferred for numerical stability (since $v$ blows up as $t\to 1$). A simple version is to modify the conditional probability path to
$p_{t}(\cdot|X=x_{1})=\mathcal{N}(tx_{1},(1-(1-\sigma_{min})t)^{2}I_{d}),$
for some small $\sigma_{min}>0$, which corresponds to the regularized conditional velocity field $v(t,z|X=x_{1})=\frac{x_{1}-(1-\sigma_{min})z}{1-(1-\sigma_{min})t}$. Another version is to consider a smoothed version of the data distribution $p^{*}$; e.g., $p_{1}=p^{*}\star\mathcal{N}(0,\sigma_{min}^{2}I_{d})$, where $\star$ denotes convolution.
@@ 3 Empirical Flow Matching
where $w_{i}(t,z)=\text{softmax}_{i}\left(\left(-\frac{1}{2(1-t)^{2}}\|z-tx^{(j)}\|^{2}\right)_{j\in[N]}\right)$, with $\text{softmax}_{i}$ denoting the $i$th component of the vector obtained after applying the softmax operation. This optimal velocity field is thus a time-dependent weighted average of the $N$ different directions towards the $x^{(i)}$. Similar formula can also be obtained for regularized versions of rectified flow.
@@ 4 FM Through the Lens of Energetics
Hence, the optimization problem (Benamou-Brenier formulation) can be written in its Eulerian form, and minimizes the total kinetic energy over all admissible paths:
@@
$\frac{\partial\hat{p}_{t}}{\partial t}+\nabla\cdot(\hat{p}_{t}\,\hat{v}(t,\cdot))=0.$ (19)
In particular, the empirical minimizer satisfies $\hat{v}^{*}(t,z)=\hat{v}(t,z)$ pointwise, and hence the pair $(\hat{p}_{t},\hat{v}^{*})$ also satisfies (19).
@@
Let the empirical target distribution be $\hat{p}_{1}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x^{(i)}}$. Consider the family of empirical affine flows defined by the conditional probability paths $p_{t}(z|x^{(i)})$ and their corresponding conditional velocity fields $v_{i}(t,z):=v(t,z|x^{(i)})=a_{t}(x^{(i)})z+b_{t}(x^{(i)})$ from Proposition 1. The minimizer $\hat{v}^{*}(t,z)$ of the empirical FM/CFM objective is a gradient field if and only if the following data-dependent condition holds for all $t\in[0,1]$:
@@
Neural network trained CFM models are implicitly optimizing for a sampler that is not only energetically inefficient (not minimizing the kinetic energy), but may also lead to memorization (producing samples that are close enough to those from the training set).
@@
It turns out that we can also obtain a similar bound for the empirical RF model $\hat{v}^{*}(t,z)$ (which is nonlinear in $z$) under the same assumption for $p_{1}$, despite the fact that it generally does not give rise to an OT map.
@@
Combining these two interpretations, we see that neural network trained CFM models are implicitly optimizing for a sampler that is not OT-optimal in expected kinetic energy (using more energy on average than the OT solution due to lack of gradient structure), and may also lead to memorization. However, regardless of optimality, both OT and CFM models generate samples with exponentially concentrated kinetic energy around typical values, with high-energy samples (which are rare) predominantly found in low-density tail regions.
@@ 5 Conclusion
In these notes, we show that empirical FM optimizes for velocity fields that generally lack a gradient structure and therefore cannot reproduce OT maps or minimize kinetic energy. This partially explains both the energetic inefficiency and the memorization tendencies observed in practice. Despite this, the generated kinetic energy concentrates sharply: exponentially for Gaussian sources and polynomially for heavy-tailed ones. This shows that the tail behavior of FM samplers is governed mainly by the source distribution, not the underlying data. Understanding these structural biases clarifies the limitations of empirical FM and may guide improved sampler designs, which we leave for future work.
@@ Appendix A Related Work
Flow Matching and Related Models. Flow Matching (FM) [lipman2022flow, lipman2024flow] and Conditional Flow Matching (CFM) [tong2023improving] have been developed as scalable alternatives [esser2024scaling] to diffusion-based generative models [song2020score, lai2025principles]. Recent work has analyzed their statistical, geometric, and algorithmic foundations, including distributional properties of FM [kunkel2025distribution], particle and bridge-based interpretations [bamberger2025carr], and geometric structure and gauge freedom in learned flows [wanelucidating, horvat2024gauge]. Extensions include guided generation [feng2025guidance], statistical efficiency analyses [mena2025statistical], rigorous comparisons between FM and optimal transport [hertrich2025relation, wald2025flow], and related studies on spatio-temporal physical systems [lim2024elucidating, erichson2025flex]. The kinetic behavior of flow-based samplers has also been examined in [shaul2023kinetic, li2025enfopath].
Memorization vs. Generalization in Generative Models. A large body of recent work studies memorization, generalization, and interpolation phenomena in modern generative models. For diffusion models, prior work has analyzed identifiability, overfitting, and deterministic sampling behavior [pidstrigach2022score, yoon2023deterministic]. Further studies provide theoretical and empirical characterizations of interpolation, dataset coverage, and memorization tendencies [lyu2025resolving, scarvelis2023closed, baptista2025memorization, bertrand2025closed, chen2025interpolation]. Broader perspectives on generalization in flow-based and likelihood-based models appear in [zhang2025understanding, gao2024flow, li2024good, ye2025provable, reu2025gradient, li2025generation]. Our work contributes to this line by showing that empirical FM induces a structural bias—non-conservative velocities—that may lead to memorization-like behavior even without model approximation error.
@@ B.3 Proof of Proposition 3
We shall apply the integral formula: for $A>0$,
$\int_{-\infty}^{\infty}e^{-Ax^{2}+Bx}\,dx=\sqrt{\frac{\pi}{A}}\exp\left(\frac{B^{2}}{4A}\right).$
The expectation is:
$\mathbb{E}[e^{aW+bW^{2}}]=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}\exp\left(-\left(\frac{1}{2}-b\right)w^{2}+aw\right)\,dw.$
Identify $A=\frac{1}{2}-b$ (which is positive since $b<1/2$) and $B=a$, and apply the formula gives:
$\mathbb{E}\left[e^{aW+bW^{2}}\right]$ $=\frac{1}{\sqrt{2\pi}}\cdot\sqrt{\frac{\pi}{\frac{1}{2}-b}}\cdot\exp\left(\frac{a^{2}}{4(\frac{1}{2}-b)}\right)=\frac{1}{\sqrt{1-2b}}\exp\left(\frac{a^{2}}{2(1-2b)}\right).$
@@
Substituting this into the energy definition, we have:
$E(Y)=\|Y-R^{-1}(Y)\|^{2}.$
Since $R^{-1}(R(X))=X$ by definition of the inverse, this simplifies to $E=\|R(X)-X\|^{2}.$ Using the definition of $R(X)$:
@@
Let $A=\Sigma_{1}^{1/2}-I_{d}$. Note that $A$ is symmetric and we can consider the eigen-decomposition $A=UDU^{T}$, where $U$ is orthogonal and $D$ is diagonal with elements $d_{i}$. The eigenvalues of $\Sigma_{1}^{1/2}$ are $\sqrt{\lambda_{i}(\Sigma_{1})}$. Thus, the eigenvalues of $A$ are $d_{i}=\sqrt{\lambda_{i}(\Sigma_{1})}-1.$ The kinetic energy can then be written as:
@@ B.5 Proof of Theorem 2
For $u\geq U_{t}$ large enough so that $s^{2}=\frac{u}{C_{K}(T)}-1\geq\frac{u}{2C_{K}(T)}$, we have:
@@
where $C_{\alpha}^{\prime}:=C_{\alpha}(2C_{K}(T))^{\alpha/2}$. This proves the first inequality in Theorem 2 with $\gamma=\alpha/2$.